OpenStudy (anonymous):

Integral of x^n/(1+x^n) is?

OpenStudy (anonymous):

there is no closed form for this integral in terms of elementary functions

OpenStudy (anonymous):

the answer is \[\int\limits_{}^{} \prime f(x)/f(x) = f(x)^{n+1}/n+1 = (1+x ^{n})^{2}/2\]